# The reflection complexity of sequences over finite alphabets

### v1

@@ Abstract
In combinatorics on words, the well-studied factor complexity function $\rho_{\mathbf{x}}$ of a sequence $\mathbf{x}$ over a finite alphabet counts, for any nonnegative integer $n$, the number of distinct length-$n$ factors of $\mathbf{x}$. In this paper, we introduce the reflection complexity function $r_{\mathbf{x}}$ to enumerate the factors occurring in a sequence $\mathbf{x}$, up to reversing the order of symbols in a word. We introduce and prove results on $r_{\mathbf{x}}$ regarding its growth properties and relationship with other complexity functions. We prove that if $\mathbf{x}$ is $k$-automatic, then $r_{\mathbf{x}}$ is computably $k$-regular, and we use the software Walnut to evaluate the reflection complexity of automatic sequences, such as the Thue–Morse sequence. We prove a Morse–Hedlund-type result characterizing eventually periodic sequences in terms of their reflection complexity, and we deduce a characterization of Sturmian sequences. Furthermore, we investigate the reflection complexity of episturmian, $(s+1)$-dimensional billiard, and Rote sequences. There are still many unanswered questions about this measure.
Keywords: reflection complexity, factor complexity, reversal, automatic sequence, Sturmian sequence, episturmian sequence, Rote sequence, Morse-Hedlund theorem, Walnut.
MSC: 05A05, 11B85, 68R15
@@ 1 Introduction
The discipline of combinatorics on words continues to be burgeoning as a relatively new and interdisciplinary area of mathematics. In this regard, the significance of combinatorics on words within disciplines such as theoretical computer science leads us to explore variants and generalizations of fundamental objects and constructions involved within the field. If $\mathbf{x}$ is an infinite sequence over a finite alphabet (see Section 1.1 for precise definitions), natural problems that arise in the combinatorial study of $\mathbf{x}$ and in the context of computer science-based problems concern the behavior of factors of $\mathbf{x}$. Writing =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,\ldots\}$, we are led to consider the factor complexity function $\rho_{\mathbf{x}}\colon\mathbb{N}_{0}\to\mathbb{N}$ which maps $n\geq 0$ to the number of distinct factors of $\mathbf{x}$ of length $n$. (The term factor refers to any contiguous block occurring in $\mathbf{x}$.)
Variations on this definition can be considered as a measure of how “complicated” a sequence is. For example, the abelian complexity function of $\mathbf{x}$ counts the number of factors of $\mathbf{x}$ of a given length and up to permutation of symbols, so that two factors $u=u(1)u(2)\cdots u(m)$ and $v=v(1)v(2)\cdots v(n)$ are abelian equivalent if $m=n$ and if there exists a bijection $\sigma\colon\{1,2,\ldots,m\}\to\{1,2,\ldots,m\}$ such that $u(1)u(2)\cdots u(m)=v(\sigma(1))v(\sigma(2))\cdots v(\sigma(m))$. Similarly, the cyclic complexity function $c_{\mathbf{x}}$ introduced in 2017 [29] is equal to the number of length-$n$ factors of $\mathbf{x}$, up to equivalence under cyclic permutations. The abelian and cyclic complexity functions lead us to introduce a reflection complexity function on sequences via an invariant property involving permutations of symbols, by analogy with the abelian and cyclic complexity functions.
In addition to the factor, abelian, and cyclic complexity functions indicated above, there have been many different complexity functions on sequences that have been previously introduced. In this regard, we highlight the following “complexities”: additive complexity [9], arithmetical complexity [12], gapped binomial complexity [77], $k$-abelian complexity [55], $k$-binomial complexity [76], Kolmogorov complexity [56], Lempel–Ziv complexity [58], Lie complexity [17], linear complexity (see the survey by Niederreiter [69]), maximal pattern complexity [54], maximum order complexity [44], (initial) (non-)repetitive complexity [25, 65], opacity complexity [8], open and closed complexity [71], palindrome complexity [4], periodicity complexity [64], privileged complexity [73], relational factor complexity [32], span and leftmost complexity [30], string attractor profile complexity (implicitly defined in [81]; also see [30]), and window complexity [31]. Also see the references in the surveys in [1, 46, 47, 48]. The complexity function $r_{\mathbf{x}}$ defined below does not seem to have been previously studied, but may be thought of as natural in terms of its relationships with automatic sequences such as the Thue–Morse sequence. To begin with, we require the equivalence relation $\sim_{r}$ defined below.
@@
Let $m,n$ be nonnegative integers. Given a finite word $u=u(1)u(2)\cdots u(m)$, its reversal is the word $u^{R}=u(m)u(m-1)\cdots u(1)$, that is, $u^{R}(i)=u(m+1-i)$ for all $i\in\{1,\ldots,m\}$. A palindrome is a word that is equal to its reversal. Two finite words $u$ and $v$ are reflectively equivalent if $u=v$ or $u=v^{R}$. We denote this equivalence relation by $u\sim_{r}v$.
@@
denote the Thue–Morse sequence, where the $n$th term in (1) for $n\geq 1$ is defined as the number of ones, modulo $2$, in the base-2 expansion of $n$. The initial terms of the integer sequence $(r_{\mathbf{t}}(n))_{n\geq 0}$ are such that
@@
The integer sequence in (2) is not currently included in the On-Line Encyclopedia of Integer Sequences [85], which suggests that our notion of “reflection complexity” is new. See also the work of Krawchuk and Rampersad in [57], which introduced the notion of cyclic/reversal complexity for sequences. The evaluation of reflection complexity functions is closely related to the work of Rampersad and Shallit [74], who investigated sequences $\mathbf{x}$ such that all sufficiently long factors $w$ have the property that that $w^{R}$ is not a factor of $\mathbf{x}$. Also, the evaluation of reflection complexities for sequences is related to the enumeration of palindromes contained in sequences; see, for example, Fici and Zamboni [49].
@@
This paper is organized as follows. In Section 1.1, we introduce the notation and definitions needed for the paper. In Section 2, we give general results on the reflection complexity. In particular, we investigate growth properties and relationships with other complexity functions. In Sections 3, 4, and 5, respectively, we investigate reflection complexities for eventually periodic sequences, Sturmian sequences and generalizations, and reversal-closed and rich sequences. Next, in Section 6, we focus on classical automatic sequences and, with the use of Walnut, we prove that the reflection complexity function for automatic sequences is a regular sequence. We also study the reflection complexity for famous automatic sequences such as the Thue–Morse sequence. Finally, some further research directions are considered in Section 7.
1.1 Preliminaries
@@
Generalities. For a general reference on words, we cite [59]. An alphabet is a finite set of elements called letters. A word on an alphabet $A$ is a finite sequence of letters from $A$. The length of a word, denoted between vertical bars, is the number of its letters (counting multiplicities). The empty word is the only 00-length word, denoted as $\varepsilon$. For all $n\geq 0$, we let $A^{n}$ denote the set of all length-$n$ words over $A$. We let $A^{*}$ denote the set of words over $A$, including the empty word and equipped with the concatenation operation. In order to distinguish finite words and infinite sequences, we write the latter in bold. Except for complexity functions, we start indexing words and sequences at $1$. A factor of a word or a sequence is any of its (finite and contiguous) subblocks. A prefix (resp., suffix) is any starting (resp., ending) factor. Given a word $w$, its $n$th term is written $w(n)$ for $1\leq n\leq|w|$. The factor starting at position $n$ and ending at position $m$ with $1\leq m\leq n\leq|w|$ is written $w[m..n]$. A factor $u$ of a word $w$ over $A$ is right (resp., left) special if $ua$ and $ub$ (resp., $au$ and $bu$) are factors of $w$ for some distinct letters $a,b\in A$. We extend these notations to (finite) words as well. A sequence $\mathbf{x}$ is reversal-closed if, for any factor $w$ of $\mathbf{x}$, the word $w^{R}$ is also a factor of $\mathbf{x}$. A sequence $\mathbf{x}$ is eventually periodic if there exist finite words $u,v$, with $v$ nonempty, such that $\mathbf{x}=uv^{\omega}$ where $v^{\omega}=vvv\cdots$ denotes the infinite concatenation of the word $v$. A sequence that is not eventually periodic is said to be aperiodic. A sequence is said to be recurrent if every factor occurs infinitely many times; it is uniformly recurrent if each factor occurs with bounded gaps, i.e., for all factors $w$, there is some length $m=m(w)$ such that $w$ occurs in every length-$m$ block.
@@
Morphisms. Let $A$ and $B$ be finite alphabets. A morphism $f\colon A^{*}\to B^{*}$ is a map satisfying $f(uv)=f(u)f(v)$ for all $u,v\in A^{*}$. In particular, $f(\varepsilon)=\varepsilon$, and $f$ is entirely determined by the images of the letters in $A$. For an integer $k\geq 2$, a morphism is $k$-uniform if it maps each letter to a length-$k$ word. A $1$-uniform morphism is called a coding. A sequence $\mathbf{x}$ is morphic if there exist a morphism $f\colon A^{*}\to A^{*}$, a coding $g\colon A^{*}\to B^{*}$, and a letter $a\in A$ such that $\mathbf{x}=g(f^{\omega}(a))$, where $f^{\omega}(a)=\lim_{n\to\infty}f^{n}(a)$. We let $E\colon\{0,1\}^{*}\to\{0,1\}^{*}$ be the exchange morphism defined by $E(0)=1$ and $E(1)=0$. We naturally extend $E$ on sequences.
@@
Numeration systems. Let $U=(U(n))_{n\geq 0}$ be an increasing sequence of integers with $U(0)=1$. Any integer $n$ can be decomposed, not necessarily uniquely, as $n=\sum_{i=0}^{t}c(i)\,U(i)$ with non-negative integer coefficients $c(i)$. The word $c(t)\cdots c(0)\in\mathbb{N}^{*}$ is a $U$-representation of $n$. If this representation is computed greedily, then for all $j<t$ we have $\sum_{i=0}^{j}c(i)\,U(i)<U(j+1)$ and $\operatorname{rep}_{U}(n)=c(t)\cdots c(0)$ is said to be the greedy $U$-representation of $n$. By convention, the greedy representation of 00 is the empty word $\varepsilon$, and the greedy representation of $n>0$ starts with a non-zero digit. For any $c_{t}\cdots c_{0}\in\mathbb{N}^{*}$, we let $\operatorname{val}_{U}(c(t)\cdots c(0))$ denote the integer $\sum_{i=0}^{t}c(i)\,U(i)$. A sequence $U$ satisfying all the above conditions defines a positional numeration system.
@@
A generalization of automatic sequences to infinite alphabets is the following [7, 75, 82]. Let $U=(U(n))_{n\geq 0}$ be an positional numeration system. A sequence $\mathbf{x}$ is $U$-regular if there exist a column vector $\lambda$, a row vector $\gamma$ and matrix-valued morphism $\mu$ such that $\mathbf{x}(n)=\lambda\mu(\operatorname{rep}_{U}(n))\gamma$. Such a system of matrices forms a linear representation of $\mathbf{x}$.
Another definition of $k$-regular sequences is the following one [7]. Consider a sequence $\mathbf{x}$ and an integer $k\geq 2$. The $k$-kernel of $\mathbf{x}$ is the set of subsequences of the form $(\mathbf{x}(k^{e}n+r))_{n\geq 0}$ where $r\in\{1,2,\ldots,k^{e}\}$. Equivalently, a sequence is $k$-regular if the $\mathbb{Z}$-module generated by its $k$-kernel is finitely generated. A sequence is then $k$-automatic if and only if its $k$-kernel is finite [7].
@@
This theorem implies in particular that $\rho_{\mathbf{x}}(n)$ is either bounded or it satisfies $\rho_{\mathbf{x}}(n)\geq n+\ell$ for all $n$. Thus the minimal factor complexity among all non-eventually periodic sequences is $\rho_{\mathbf{x}}(n)=n+1$ for all $n$, i.e., the complexity of Sturmian sequences. Actually there is another “growth gap” for $\rho_{\mathbf{x}}$. Recall that a sequence $\mathbf{x}$ is called quasi-Sturmian if there exists a constant $C$ such that, for $n$ large enough, one has $\rho_{\mathbf{x}}(n)=n+C$ (see [28]; also see [34]). It is known that if $\mathbf{x}$ is neither eventually periodic nor quasi-Sturmian, then $\rho_{\mathbf{x}}(n)-n$ tends to infinity (this result is due to Coven [37, Lemma 1.3]; the first author [3, Theorem 3, p. 23] attributed the result to J. Cassaigne, because he first learned it from him). Thus we have:
@@
One more point (once explained to the first author by J. Berstel) is that if $\rho_{\mathbf{x}}(n)=n+1$ for $n$ large enough, then $\rho_{\mathbf{x}}(n)=n+1$ for all $n$. Namely, let $n_{0}$ be the least integer $n$ for which $\rho_{\mathbf{x}}(n)=n+1$, and suppose that $n_{0}>1$. Hence $\rho_{\mathbf{x}}(n_{0}-1)\neq n_{0}$. The sequence $\mathbf{x}$ cannot be eventually periodic, since its factor complexity is not bounded. Thus, one has $\rho_{\mathbf{x}}(n)\geq n+1$ for all $n$. Hence, in particular, $\rho_{\mathbf{x}}(n_{0}-1)\geq n_{0}$. Thus $\rho_{\mathbf{x}}(n_{0}-1)>n_{0}$. Since $\rho_{\mathbf{x}}$ is non-decreasing, we have that $n_{0}<\rho_{\mathbf{x}}(n_{0}-1)\leq\rho_{\mathbf{x}}(n_{0})=n_{0}+1$. This gives $\rho_{\mathbf{x}}(n_{0}-1)=n_{0}+1=\rho_{\mathbf{x}}(n_{0})$, which is impossible since $\mathbf{x}$ is not eventually periodic. In other words, in the second item above, if $C=1$, then $\mathbf{x}$ is Sturmian.

### v4

@@ Abstract
In combinatorics on words, the well-studied factor complexity function $\rho_{\mathbf{x}}$ of a sequence $\mathbf{x}$ over a finite alphabet counts, for every nonnegative integer $n$, the number of distinct length-$n$ factors of $\mathbf{x}$. In this paper, we introduce the reflection complexity function $r_{\mathbf{x}}$ to enumerate the factors occurring in a sequence $\mathbf{x}$, up to reversing the order of symbols in a word. We prove a number of results about the growth properties of $r_{\mathbf{x}}$ and its relationship with other complexity functions. We also prove a Morse–Hedlund-type result characterizing eventually periodic sequences in terms of their reflection complexity, and we deduce a characterization of Sturmian sequences. We investigate the reflection complexity of quasi-Sturmian, episturmian, $(s+1)$-dimensional billiard, complementation-symmetric Rote, and rich sequences. Furthermore, we prove that if $\mathbf{x}$ is $k$-automatic, then $r_{\mathbf{x}}$ is computably $k$-regular, and we use the software Walnut to evaluate the reflection complexity of some automatic sequences, such as the Thue–Morse sequence. We note that there are still many unanswered questions about this reflection measure.
Jean-Paul Allouche    John M. Campbell    Shuo Li
Jeffrey Shallit    Manon Stipulanti
Keywords: factor complexity, reflection complexity, reversal, automatic sequence, Sturmian sequence, quasi-Sturmian sequence, episturmian sequence, billiard sequence, Rote sequence, rich sequence, Morse-Hedlund theorem, Walnut.
MSC: Primary 68R15; Secondary 11B85.
@@ 1 Introduction
Given an infinite sequence $\mathbf{x}$ over a finite alphabet, it is natural to study the combinatorial properties of the factors of $\mathbf{x}$. (The term factor refers to a contiguous block occurring in $\mathbf{x}$.) For example, writing $\mathbb{N}_{0}=\{0,1,\ldots\}$ and =$\mathbb{N}=\{1,2,\ldots\}$, many authors have studied the factor complexity function $\rho_{\mathbf{x}}\colon\mathbb{N}_{0}\to\mathbb{N}$, which maps $n\geq 0$ to the number of distinct factors of $\mathbf{x}$ of length $n$. Note that $\rho_{\mathbf{x}}(0)=1$, since every—finite or infinite—sequence has a unique factor of length 00; namely, the empty word.
Variations on this definition can be considered as a measure of how “complicated” a sequence is. For example, the abelian complexity function of $\mathbf{x}$ counts the number of factors of $\mathbf{x}$ of a given length, where two factors $u$ and $v$ are considered the same if they have the same length and one is a permutation of the other. Similarly, the cyclic complexity function $c_{\mathbf{x}}$, introduced in 2017 [32], is equal to the number of length-$n$ factors of $\mathbf{x}$, up to equivalence under rotations (cyclic permutations). By analogy, the abelian and cyclic complexity functions lead us to introduce, in this paper, a reflection complexity function on sequences involving reversals.
In addition to the factor, abelian, and cyclic complexity functions indicated above, there have been many different complexity functions on sequences that have been previously introduced. In this regard, we highlight the following in alphabetical order: additive complexity [9], arithmetical complexity [12], gapped binomial complexity [84], $k$-abelian complexity [60], $k$-binomial complexity [83], Kolmogorov complexity [61], Lempel–Ziv complexity [63], Lie complexity [17], linear complexity [75], maximal pattern complexity [59], maximum order complexity [48], opacity complexity [8], open and closed complexity [77], palindrome complexity [4], periodicity complexity [70], privileged complexity [79], relational factor complexity [35], (initial) (non-)repetitive complexity [28, 71], span and leftmost complexity [33], string attractor profile complexity (implicitly defined in [88]; also see [33]), and window complexity [34]. Also see the references in the surveys in [1, 50, 51, 52]. Our reflection complexity function $r_{\mathbf{x}}$, defined below, does not seem to have been previously studied, but may be thought of as natural in terms of its relationships with automatic sequences such as the Thue–Morse sequence. To begin with, we require the equivalence relation $\sim_{r}$ defined below.
@@
Let $m,n$ be nonnegative integers. Given a finite word $u=u(1)u(2)\cdots u(m)$, its reversal is the word $u^{R}=u(m)u(m-1)\cdots u(1)$, i.e., $u^{R}(i)=u(m+1-i)$ for all $i\in\{1,\ldots,m\}$. A palindrome is a word that is equal to its reversal. Two finite words $u$ and $v$ are reflectively equivalent if $v=u$ or $v=u^{R}$. We denote this equivalence relation by $u\sim_{r}v$.
@@
denote the Thue–Morse sequence, where the $n$th term in (1) for $n\geq 1$ is defined as the number of $1$’s, modulo $2$, in the base-2 expansion of $n-1$. The initial terms of the integer sequence $(r_{\mathbf{t}}(n))_{n\geq 0}$ are such that
@@
The integer sequence in (2) was not, prior to this paper, included in the On-Line Encyclopedia of Integer Sequences [92], which suggests that our notion of “reflection complexity” is new. (Now it is present as sequence A373700.) Also see the work of Krawchuk and Rampersad in [62], which introduced the notion of cyclic/reversal complexity for sequences. The evaluation of reflection complexity functions is closely related to the work of Rampersad and Shallit [80], who investigated sequences $\mathbf{x}$ such that all sufficiently long factors $w$ have the property that $w^{R}$ is not a factor of $\mathbf{x}$. Also, the evaluation of reflection complexities for sequences is related to the enumeration of palindromes contained in sequences; see, e.g., Fici and Zamboni [53].
@@
This paper is organized as follows. In Section 2, we introduce the notation and definitions needed for the paper. In Section 3, we give general results about reflection complexity. In particular, we investigate its growth properties and relationships with other complexity functions. In Section 4, we give a graph-theoretic interpretation of reflection-equivalent classes and prove an inequality for reflection complexity. In Sections 5, 6, and 7, respectively, we investigate reflection complexity for eventually periodic sequences, Sturmian sequences and generalizations, and reversal-closed and rich sequences. Next, in Section 8, we focus on classical automatic sequences and, with the use of the free software Walnut, we prove that the reflection complexity function for automatic sequences is a regular sequence. We also study reflection complexity for famous automatic sequences such as the Thue–Morse sequence. Finally, some further research directions and open questions are considered in Section 9.
2 Preliminaries
@@
Generalities. For a general reference on words, we cite [65]. An alphabet is a finite set of elements called letters. A word over an alphabet $A$ is a finite sequence of letters from $A$. The length of a word, denoted between vertical bars, is the number of its letters (counting multiplicities). The empty word is the only 00-length word, denoted by $\varepsilon$. For all $n\geq 0$, we let $A^{n}$ denote the set of all length-$n$ words over $A$. We let $A^{*}$ denote the set of words over $A$, including the empty word and equipped with the concatenation operation. In order to distinguish finite words and infinite sequences, we write the latter in bold. Except for complexity functions, we start indexing words and sequences at $1$, unless otherwise specified. A factor of a word or a sequence is one of its (finite and contiguous) subblocks. A prefix (resp., suffix) is a starting (resp., ending) factor. Given a word $w$, its $n$th term is written $w(n)$ for $1\leq n\leq|w|$. The factor starting at position $n$ and ending at position $m$ with $1\leq m\leq n\leq|w|$ is written $w[m..n]$. We let $\operatorname{Fac}_{w}$ denote the set of all factors of $w$ and, for each natural number $n$, we let $\operatorname{Fac}_{w}(n)$ denote the set of all length-$n$ factors of $w$. A factor $u$ of a word $w$ over $A$ is right (resp., left) special if $ua$ and $ub$ (resp., $au$ and $bu$) are factors of $w$ for some distinct letters $a,b\in A$. A sequence $\mathbf{x}$ is reversal-closed if, for every factor $w$ of $\mathbf{x}$, the word $w^{R}$ is also a factor of $\mathbf{x}$. A sequence $\mathbf{x}$ is eventually periodic if there exist finite words $u,v$, with $v$ nonempty, such that $\mathbf{x}=uv^{\omega}$ where $v^{\omega}=vvv\cdots$ denotes the infinite concatenation of $v$. A sequence that is not eventually periodic is said to be aperiodic. A sequence is said to be recurrent if every factor occurs infinitely many times; it is uniformly recurrent if each factor occurs with bounded gaps, i.e., for all factors $w$, there is some length $m=m(w)$ such that $w$ occurs in every length-$m$ block.
@@
Morphisms. Let $A$ and $B$ be finite alphabets. A morphism $f\colon A^{*}\to B^{*}$ is a map satisfying $f(uv)=f(u)f(v)$ for all $u,v\in A^{*}$. In particular, $f(\varepsilon)=\varepsilon$, and $f$ is entirely determined by the images of the letters in $A$. For an integer $k\geq 2$, a morphism is $k$-uniform if it maps each letter to a length-$k$ word. A $1$-uniform morphism is called a coding. A sequence $\mathbf{x}$ is pure morphic if it is a fixed point of a morphism, i.e., there exist a morphism $f\colon A^{*}\to A^{*}$ and a letter $a\in A$ such that $\mathbf{x}=f^{\omega}(a)$, where $f^{\omega}(a)=\lim_{n\to\infty}f^{n}(a)$. A sequence $\mathbf{y}$ is morphic if there exist a pure morphic sequence $\mathbf{x}$ and a coding $g\colon A^{*}\to B^{*}$ such that $\mathbf{y}=g(\mathbf{x})$. We let $E\colon\{0,1\}^{*}\to\{0,1\}^{*}$ be the exchange morphism defined by $E(0)=1$ and $E(1)=0$. We naturally extend $E$ to sequences.
@@
Numeration systems. Let $U=(U(n))_{n\geq 0}$ be an increasing sequence of integers with $U(0)=1$. Any integer $n$ can be decomposed in a greedy way as $n=\sum_{i=0}^{t}c(i)\,U(i)$ with non-negative integer coefficients $c(i)$. The word $c(t)\cdots c(0)\in\mathbb{N}^{*}$ is said to be the (greedy) $U$-representation of $n$. By convention, the greedy representation of 00 is the empty word $\varepsilon$, and the greedy representation of $n>0$ starts with a non-zero digit. For $c_{t}\cdots c_{0}\in\mathbb{N}^{*}$, we let $\operatorname{val}_{U}(c(t)\cdots c(0))$ denote the integer $\sum_{i=0}^{t}c(i)\,U(i)$. A sequence $U$ satisfying all the above conditions defines a positional numeration system.
@@
A generalization of automatic sequences to infinite alphabets is the following [7, 82, 89]. Let $U=(U(n))_{n\geq 0}$ be a positional numeration system. A sequence $\mathbf{x}$ is $U$-regular if there exist a column vector $\lambda$, a row vector $\gamma$ and matrix-valued morphism $\mu$ such that $\mathbf{x}(n)=\lambda\mu(\operatorname{rep}_{U}(n))\gamma$. Such a system of matrices forms a linear representation of $\mathbf{x}$. In particular, if $U$ is the sequence of consecutive powers of an integer $k\geq 2$, then $\mathbf{x}$ is said to be $k$-regular.
Another definition of $k$-regular sequences is the following one [7]. Consider a sequence $\mathbf{x}$ and an integer $k\geq 2$. The $k$-kernel of $\mathbf{x}$ is the set of subsequences of the form $(\mathbf{x}(k^{e}n+r))_{n\geq 0}$ where $r\in\{1,2,\ldots,k^{e}\}$. A sequence is $k$-regular if the $\mathbb{Z}$-module generated by its $k$-kernel is finitely generated. A sequence is then $k$-automatic if and only if its $k$-kernel is finite [7].
@@
This theorem implies in particular that $\rho_{\mathbf{x}}(n)$ is either bounded or it satisfies $\rho_{\mathbf{x}}(n)\geq n+\ell-1$ for all $n$. Thus the minimal factor complexity among all non-eventually periodic sequences is $\rho_{\mathbf{x}}(n)=n+1$ for all $n$, i.e., the complexity of Sturmian sequences. Actually there is another “growth gap” for $\rho_{\mathbf{x}}$. Recall that a sequence $\mathbf{x}$ is called quasi-Sturmian if there exists a constant $C$ such that, for $n$ large enough, one has $\rho_{\mathbf{x}}(n)=n+C$ (see [31]; also see [37]). It is known that if $\mathbf{x}$ is neither eventually periodic nor quasi-Sturmian, then $\rho_{\mathbf{x}}(n)-n$ tends to infinity: this result is due to Coven [40, Lemma 1.3]; also see Cassaigne’s proof cited in [3, Proof of Theorem 3, p. 23]. Thus
@@
One more point (once explained to the first author by Jean Berstel) is that if $\rho_{\mathbf{x}}(n)=n+1$ for $n$ large enough, then $\rho_{\mathbf{x}}(n)=n+1$ for all $n$. Namely, let $n_{0}$ be the least integer $n$ for which $\rho_{\mathbf{x}}(n)=n+1$, and suppose that $n_{0}>1$. Hence $\rho_{\mathbf{x}}(n_{0}-1)\neq n_{0}$. The sequence $\mathbf{x}$ cannot be eventually periodic, since its factor complexity is not bounded. Thus, one has $\rho_{\mathbf{x}}(n)\geq n+1$ for all $n$. Hence, in particular, $\rho_{\mathbf{x}}(n_{0}-1)\geq n_{0}$. Thus $\rho_{\mathbf{x}}(n_{0}-1)>n_{0}$. Since $\rho_{\mathbf{x}}$ is non-decreasing, we have that $n_{0}<\rho_{\mathbf{x}}(n_{0}-1)\leq\rho_{\mathbf{x}}(n_{0})=n_{0}+1$. This gives $\rho_{\mathbf{x}}(n_{0}-1)=n_{0}+1=\rho_{\mathbf{x}}(n_{0})$, which is impossible since $\mathbf{x}$ is not eventually periodic. In other words, in the second item above, if $C=1$, then $\mathbf{x}$ is Sturmian.